OpenStudy (anonymous):

Melinda had $2600 and April had $1500 when they went to Bali. Between then they spent $2900. If Melinda finished up with twice as much as Apirl, how much did April spend?

OpenStudy (anonymous):

April spent $966.6 and Melinda spent 2*966.6 = $1933.2

OpenStudy (anonymous):

sorry, i don't think you got the right answer

OpenStudy (anonymous):

Ok, can you give a hint

OpenStudy (anonymous):

i got the answer but i don't know how to solve it. April spent $1100.

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

According to the question Melinda spends x and Apirl spends y, so we can eqate that x + y = 2900 & 2600-x = 2(1500-y) since Melinda finished up with twice as much as Apirl 2600 - x = 3000 - 2y 2y - x = 400 2y - 2900 + y = 400 3y = 3300 y = 1100 and x = 2200

OpenStudy (anonymous):

if melinda had 2600 and april 1500 that equals 4100. they spent 2900 4100-2900=1200 melinda now had 800 and april now has 400

OpenStudy (anonymous):

X + Y = is meant to be 2900 not 3300.

OpenStudy (anonymous):

Therefore you are not quite right..........

OpenStudy (anonymous):

I think its right please check it once again

OpenStudy (anonymous):

yeah you're right, finding another solution

OpenStudy (anonymous):

I think Nannyx11 got a good and simple way to solve this problem .

OpenStudy (anonymous):

yes, right and btw x = 1800

OpenStudy (anonymous):

thanks any way